# Gap solitons of the Wannier and Bloch types in spin-orbit-coupled Bose-Einstein condensates with a moiré lattice

### v1

@@ IV Conclusions
We have systematically investigated GSs (gap solitons) in the dynamical models of the spin-1 BECs, under the action of the Rashba SOC (spin-orbit coupling) and a periodic or quasiperiodic ML (moiré lattice). We focused on the GSs bifurcating from the five lowest energy bands. GSs in periodic systems are classified into two species, WT (Wannier-type) and BT (Bloch-type) ones, according to their ability (WT) or inability (BT) to maintain the localization at vanishingly small norms. When the ML depth and period are sufficiently large, the lowest five energy bands become flat, and five WT GS families bifurcate from them, each approximated by a single Wannier function. More complex WT GSs are also constructed via the coherent superposition of the fundamental WT GSs with equal chemical potentials. The WT GSs are stable only at small norms. Conversely, when the ML depth and period are small, the energy bands exhibit finite bandwidths. Guided by the bifurcation characteristics of WT GSs, we have identified BT GSs whose phases patterns are similar to those of WT GSs. The broad energy bands cause the width of BT GSs to significantly expand near band edges, thereby inducing instability. Additionally, the Rashba SOC effect has been demonstrated to enhance the soliton localization while reducing the norm threshold above which the GSs exist. These findings suggest that, even with small ML depth and period, strongly localized WT GSs can still be generated by increasing the SOC strength. We have also established that quasiperiodic MLs affect spatial profiles and existence domains of the GSs, without altering their core stability mechanisms. Edge states induced by finite boundaries in quasiperiodic MLs also suggest a promising topic for the extension of the work. Overall, our results offer a theoretical basis for relevant experimental investigations, and propose a systematic approach to studies of GSs in complex periodic and quasi-periodic systems.

### v2

@@ IV Conclusions
We have systematically investigated GSs (gap solitons) in the dynamical models of the spin-1 BECs under the action of the Rashba SOC (spin-orbit coupling) in a periodic or quasiperiodic ML (moiré lattice). We focused on the GSs bifurcating from the five lowest-energy bands. GSs in periodic systems are classified into two species, WT (Wannier-type) and BT (Bloch-type), according to their ability (WT) or inability (BT) to maintain the localization at vanishingly small norms. When the ML depth and period are sufficiently large, the lowest five energy bands become flat and five WT GS families bifurcate from them, each approximated by a single Wannier function. More complex WT GSs are also constructed via the coherent superposition of the fundamental WT GSs with equal chemical potentials. The WT GSs are stable only at small norms. Conversely, when the ML depth and period are small, the energy bands exhibit finite bandwidths. Guided by the bifurcation characteristics of WT GSs, we have identified BT GSs whose phase patterns are similar to those of WT GSs. The broad energy bands cause the width of BT GSs to significantly expand near band edges, thereby inducing instability. Additionally, the Rashba SOC effect has been demonstrated to enhance the soliton localization while reducing the norm threshold above which the GSs exist. These findings suggest that, even with small ML depth and period, strongly localized WT GSs can still be generated by increasing the SOC strength. We have also established that quasiperiodic MLs affect the spatial profiles and existence domains of the GSs without altering their core stability mechanisms. The edge states induced by finite boundaries in quasiperiodic MLs also constitute a promising topic for future investigation. Overall, our results offer a theoretical basis for relevant experimental investigations and propose a systematic approach to studies of GSs in complex periodic and quasiperiodic systems.